OpenStudy (anonymous):

A=a+b/2; Solve for b. Please explain :)

jimthompson5910 (jim_thompson5910):

Is the equation \[\Large A = \frac{a+b}{2}\] OR is it\[\Large A = a+\frac{b}{2}\] ???

OpenStudy (anonymous):

It's the first one

jimthompson5910 (jim_thompson5910):

ok thanks

jimthompson5910 (jim_thompson5910):

We want to isolate b, so we need to move everything but b over to the left side

jimthompson5910 (jim_thompson5910):

So we start off by moving that '2' over. We do this by multiplying both sides by 2. \[\Large A = \frac{a+b}{2}\] \[\Large 2A = \cancel{2}\frac{a+b}{\cancel{2}} ... \ \text{Multiply both sides by 2}\] \[\Large 2A = a+b ... \ \text{Simplify}\] Do you see where to go from here?

OpenStudy (anonymous):

Yes, I do. Thanks so much, I understand :)

jimthompson5910 (jim_thompson5910):

that's great, yw